Prove

tan(x) . Cos(x) =sin(x)


1
Expert's answer
2021-01-14T20:07:28-0500

Using a right angled triangle whose sides are hypothenuse a, opposite b,and adjacent c;

sin(X)=opposite(b)/hypothenuse (a)

tan(X)=opposite (b)/adjacent (c).

Cos(X)=adjacent (c)/hypothenuse (a).

Therefore: tan(X) •cos(X)=(b/c)×(c/a)=b/a=sin(X).